OpenStudy (anonymous):

What is the equation of the line that passes through the points (0, 5) and (2, 3)?

OpenStudy (anonymous):

@Lost??? @Leigha @maria.alison @Jonnychewy

OpenStudy (texaschic101):

first thing to do is find the slope by using the slope formula slope(m) = (y2 - y1) / (x2 - x1) (0,5)....x1 = 0 and y1 = 5 (2,3),,,x2 = 2 and y2 = 3 now we sub slope(m) = (3 - 5) / (2 - 0) slope = -2/2 slope = -1 now use y = mx + b slope(m) = -1 use either of your sets of points...(2,3)..x = 2 and y = 3 now we sub and since we already know the slope(m) we need to find the y intercept(b) 3 = -1(2) + b 3 = -2 + b 3 + 2 = b 5 = b so your equation is : y = -x + 5